### v1

# BrowseComp-V3: A Visual, Vertical, and Verifiable Benchmark for Multimodal Browsing Agents

Abstract
Multimodal large language models (MLLMs), leveraging their increasingly advancing autonomous planning and tool use capabilities, are evolving into intelligent agents capable of performing web browsing for multimodal deep search. However, existing benchmarks remain limited in terms of task complexity, information searchability, and evaluation dimensions, thereby hindering comprehensive assessments of multimodal browsing agents’ deep search capabilities in open-world environments. To bridge these gaps, we present BrowseComp-V3, a novel benchmark comprising 300 meticulously hand-crafted, challenging questions across diverse domains. By emphasizing deep, multi-level, and cross-modal multi-hop reasoning, we ensure that these tasks necessitate the use of web browsing tools and cannot be resolved solely through the model’s parametric knowledge. Moreover, we strictly enforce the public searchability of all supporting evidence and incorporate an expert-validated, subgoal-driven process evaluation mechanism, thereby enabling fine-grained characterization of search behaviors and systematic analysis of capability boundaries. Beyond the dataset, we provide OmniSeeker, a general multimodal browsing agent framework, and conduct a comprehensive evaluation on MLLMs. The results demonstrate that even state-of-the-art models, such as GPT-5.2, achieve only 36% accuracy. Further analysis reveals critical bottlenecks in existing models regarding multimodal information integration and fine-grained perception, highlighting a fundamental lack of native multimodal reasoning capabilities.
1 Introduction
Multimodal large language models OpenAI [2025b]; Pichai et al. [2025]; Bai et al. [2025]; Meta [2025]; Li et al. [2024a]; Shi et al. [2025b] have demonstrated substantial performance gains across complex tasks. By integrating linguistic comprehension, visual perception, and tool-use capabilities, these models are increasingly evolving into autonomous agents capable of independent exploration and decision-making. Consequently, an increasing body of research is exploring how MLLMs can leverage external search and browsing tools to address multimodal deepsearch challenges in open-world environments OpenAI [2025a]; Google [2024]; Wu et al. [2025b]; Geng et al. [2025]; Huang et al. [2026].
Despite the rapid evolution of model capabilities, benchmarks for multimodal browsing and deep search remain noticeably underdeveloped. Existing studies Jiang et al. [2024]; Geng et al. [2025]; Li et al. [2025b]; Tao et al. [2025a] frequently exhibit shortcomings in task complexity, information searchability, and evaluation dimensions, hindering fair, holistic, and reproducible assessments of multimodal browsing agents. Existing methods still exhibit certain limitations: i) Insufficient Task Complexity. Early benchmarks Cheng et al. [2025]; Geng et al. [2025] are predominantly confined to shallow retrieval within two hops, with visual information concentrated in the initial stage. Consequently, they fail to reflect the intricacies of real-world, deep multimodal search scenarios. ii) Inaccessibility of Key Information. The core evidence in subsequent benchmarks Fu et al. [2025b]; Li et al. [2025b]; Tao et al. [2025a] is often derived from sources that are not publicly searchable by tools, such as videos or proprietary documents, which undermines the reproducibility and fairness. iii) Narrow Evaluation Dimensions. Existing studies Jiang et al. [2024]; Li et al. [2024b] primarily focus on the accuracy of the final answer but lack a systematic characterization of the reasoning process. This makes it challenging to diagnose specific failure modes or define the capability boundaries of the models.
To address these gaps, we present BrowseComp-V3, a novel benchmark specifically designed to evaluate multimodal deep browsing and search capabilities. BrowseComp-V3 comprises 300 carefully curated, highly complex questions spanning 24 distinct sub-domains, which systematically assess multimodal browsing agents in open-world settings. A key feature of our work is the emphasis on deep, multi-level, and cross-modal reasoning, where critical evidence is strategically interleaved across textual and visual modalities within and across web pages. This design effectively precludes "shortcut" successes derived solely from text-based heuristics or models’ reliance on internal parametric knowledge. Furthermore, we ensure that all critical evidence is accessible via standard public search engines and provide manually annotated gold-standard search trajectories to guarantee fairness and reproducibility. Finally, we introduce expert-validated intermediate sub-goals for each task, enabling fine-grained evaluation of the search process to precisely identify the capability boundaries and failure modes of the evaluated models. Our primary contributions are summarized as follows:
  • We present BrowseComp-V3, which, to the best of our knowledge, represents the first multimodal deep search benchmark to concurrently feature extensive search depth, public search accessibility, and process-oriented evaluation mechanisms.
  • We systematically define and categorize multimodal deep search scenarios. Through process-oriented evaluation, we provide a more comprehensive characterization of multimodal browsing agents’ capabilities and limitations.
  • We develop OmniSeeker, a unified multimodal browsing agent framework. By integrating diverse web search and visual perception tools, OmniSeeker rivals the performance of state-of-the-art closed-source systems and substantially enhances open-source models’ performance on multimodal deep search tasks.
Table 1: Comparison of our benchmark against representative deep search benchmarks along eight dimensions. § indicates that the benchmark only partially satisfies the corresponding criterion.
Benchmarks Multimodal Context inputs Multi-round Interaction(>2) Thinking with Images Multi-image Reasoning Public-search Answerable Hop-based Difficulty Analysis Human-validated Trajectories Fine-grained Progress Metrics
InfoSeek Chen et al. [2023]
Enc-VQA Mensink et al. [2023]
MMSearch Jiang et al. [2024]
DynVQA Li et al. [2024b] §
SimpleVQA Cheng et al. [2025]
LiveVQA Fu et al. [2025b] § §
BrowseComp Wei et al. [2025]
FactualVQA Wu et al. [2025b]
BrowseComp-VL Geng et al. [2025]
MM-BrowseComp Li et al. [2025b] §
MMSearch-Plus Tao et al. [2025a] §
BrowseComp-V3 (Ours)
2 Related Work
2.1 Multimodal Large Language Models
Multimodal large language models OpenAI [2025b]; Pichai et al. [2025]; ByteDance Seed [2025]; Bai et al. [2025] have demonstrated remarkable proficiency across a diverse spectrum of tasks, such as VQA Fu et al. [2025a]; Cheng et al. [2025]; Zhang et al. [2025]; Wu et al. [2025b], grounding Kazemzadeh et al. [2014], OCR Masry et al. [2022]; Shi et al. [2025a]; Mathew et al. [2021]; Shi et al. [2025c], and multimodal reasoning Lu et al. [2023]; Wang et al. [2025, 2024]. Nevertheless, MLLMs inherently struggle with the real-time acquisition of up-to-date information, posing substantial hurdles when addressing knowledge-intensive or information-retrieval queries. Consequently, contemporary research has pivoted toward tool-augmented frameworks to empower MLLMs as autonomous agents, capable of dynamically retrieving and incorporating external knowledge.
2.2 Tool-Enhanced Browsing Agents
Driven by the escalating tool-calling proficiency of LLMs/MLLMs Guo et al. [2025]; Yang et al. [2025]; OpenAI [2025b], tool-enhanced browsing agents have emerged as a pivotal research frontier. To enable precise retrieval and reasoning in dynamic web environments, recent studies advocate leveraging supervised fine-tuning and reinforcement learning to enhance agents’ reasoning and decision-making capabilities Jin et al. [2025]; Li et al. [2025a]; Wu et al. [2025a]; Tao et al. [2025b]; Song et al. [2025]; Zheng et al. [2025]. This paradigm, initially validated in textual agents, has been rapidly extended to the multimodal domain Wu et al. [2025b]; Mei et al. [2025]; Hong et al. [2025]; Huang et al. [2026]. This evolution has significantly expanded the search depth and adaptive boundaries of agents when navigating complex tasks.
2.3 Multimodal Browsing Benchmarks
Traditional multimodal browsing benchmarks typically decouple visual understanding from text retrieval and focus on simple two-hop retrieval tasks Chen et al. [2023]; Mensink et al. [2023]; Cheng et al. [2025]; Geng et al. [2025]. As visual agents advance, performance on such tasks has largely saturated. BrowseComp Wei et al. [2025] evaluates text-only agents in open-world settings by requiring large-scale web navigation, offering valuable guidance for multimodal task construction. Inspired by this paradigm, recent benchmarks such as MM-BrowseComp Li et al. [2025b] and MMSearch-Plus Tao et al. [2025a] incorporate multi-hop designs and fine-grained visual reasoning to enhance reasoning depth. However, existing benchmarks still suffer from key limitations: critical information often resides in videos or non-searchable documents, tool support is insufficient, and evaluation primarily measures final-answer correctness while overlooking the quality of reasoning. To bridge this gap, we propose BrowseComp-V3, ensuring all critical information comes from publicly accessible resources during task design. We also introduce the Process Score metric to evaluate multimodal browsing agents comprehensively.
Figure 1: An overview of the data construction process of BrowseComp-V3.
3 The BrowseComp-V3 Dataset
BrowseComp-V3 is developed by a dedicated team of over 20 researchers, including Master’s and Ph.D. candidates with expertise in artificial intelligence and related fields. The entire workflow adheres to predefined design principles and a multi-stage quality control pipeline, as delineated in the following subsections.
3.1 Design Principles
BrowseComp-V3 follows 3 core design principles that address key limitations of existing benchmarks in task complexity, information searchability, and evaluation dimensions.
Multi-dimensional Cross-modal Coverage.
To more faithfully simulate real-world search scenarios, we augment task complexity along two distinct dimensions. Specifically, we extend search depth via multi-hop variations and categorize cross-modal interaction complexities into 3 hierarchical levels: intra-region alignment, inter-region integration, and inter-image reasoning.
Process-oriented Granular Evaluation.
Datasets should incorporate expert-validated sub-goals to enable systematic tracking of intermediate reasoning steps. This design ensures granular tracking of evidence acquisition phases, thereby permitting a rigorous diagnostic analysis of failure modes and an accurate delineation of model capability boundaries.
High Reliability and Reproducibility.
For rigorous evaluation, we adopt 3 filtering criteria: (1) Evidence Traceability. Require all evidence be publicly accessible through search tools with complete manual annotation trajectories. (2) Temporal Stability. Prioritize temporally invariant, objective knowledge to eliminate dynamic web content fluctuations. (3) Answer Objectivity. Enforce concise, verifiable answers to enable standardized automated evaluation.
3.2 Data Construction Pipeline
As illustrated in Figure 1, BrowseComp-V3  construction follows a closed-loop quality assurance framework comprising 5 stages:
Stage 1: Initialization and Guideline Formulation
The experts defines core evaluation dimensions (domain diversity, task hierarchy, and hop distribution) and constructs Initial Exemplars comprising Visual Inputs, Queries, Sub-goals, Answer, and Metadata. These exemplars, together with Instruction Documents, establish the gold standard for subsequent large-scale annotation.
Stage 2: Tool-Augmented Exploratory Annotation
Annotators are assigned sub-tasks according to domain expertise and conduct exploratory web searches using a suite of specialized tools, including TextSearch, WebVisit, ImageSearch, ImageCrop, and ReverseImageSearch. They document complete interaction trajectories, partition complex tasks into pivotal sub-goals, and annotate the capabilities required to acquire each critical piece of evidence.
Stage 3: Dual-Verification and Adversarial Filtering
The original dataset undergoes two sequential screening phases. First, in the human verification loop, verifiers replicate the annotated search trajectories and evaluate logical coherence, evidentiary support, and answer accuracy. Samples that fail verification are returned for revision. Second, state-of-the-art (SOTA) multimodal large models filter out trivial examples, ensuring the retention of challenging samples that involve long-tail knowledge or complex reasoning requirements.
Stage 4: Structured Data Formatting
The verified samples are post-processed and converted into a unified JSON format, with standardized input/output fields, sub-goals, and interaction trajectories. This formatting ensures both human readability and machine interpretability, enabling automated evaluation pipelines.
Stage 5: Expert Quality Control
Before the formal release, domain experts audit the structured data for safety, privacy compliance, and factual accuracy. Only approved samples are included in the final dataset, ensuring ethical and professional standards.
Type Statistic Number
Basic Statistics Total questions 300
Total images 383
Maximum question length 134
Maximum answer length 23
Average question length 58.58
Average answer length 2.47
Category Statistics Primary 5
Secondary 24
Task Level Level 1 89
Level 2 140
Level 3 71
Difficulty Distribution Easy 45
Medium 139
Hard 86
Expert 30
Figure 2: Statistics of BrowseComp-V3. (Left) Category distribution across primary domains. (Right) Summary of statistics.
3.3 Dataset Statistics
Figure  2 (left) illustrates the categorical distribution of  BrowseComp-V3. The dataset comprises 5 balanced categories: Science, Technology, Society, Culture, and Life. Additional statistical metrics, including basic statistics, task levels and difficulty distributions, are provided in Figure  2 (right).
4 Experiments
4.1 Experimental Setup
Evaluated Models
We systematically evaluate BrowseComp-V3under 4 representative settings, as detailed below:
  • Human. To assess human performance, we recruit participants with PhD-level expertise who independently solve each problem utilizing a standard web browser. Participants can freely browse publicly accessible web resources to gather evidence and produce verifiable answers.
  • Tool-Free MLLMs. We benchmark multiple SOTA MLLMs in a tool-free setting, where models must generate answers directly without access to external tools or search capabilities. Specifically, we evaluate the following models: GPT-5.2 OpenAI [2025b], o4-mini OpenAI [2025c], GPT-4o Hurst et al. [2024], Gemini-3-Flash-Preview Google [2025], Claude-Sonnet-4.5 Anthropic [2025], Doubao-Seed-1.8 ByteDance Seed [2025], MiMo-V2-Flash Team et al. [2026], Qwen3-VL-235B-A22B-Instruct Bai et al. [2025], and Qwen3-VL-8B-Instruct Bai et al. [2025].
  • Tool-Augmented MLLMs. Additionally, we evaluate tool-augmented model services accessed through their official web platforms, with the maximum reasoning mode enabled to elicit their full capabilities. Concretely, we evaluate the following models: GPT-5.2-Thinking OpenAI [2026], Gemini-3-Pro-Preview Google [2026], and Claude-Sonnet-4.5-Thinking Anthropic [2026].
  • OmniSeeker. Lastly, we evaluate models using OmniSeeker, our custom-built multimodal browsing agent—a unified and transparent framework equipped with standardized tools, including TextSearch, WebVisit, ImageSearch, ImageCrop, and ReverseImageSearch.
Implementation Details
We employ a unified and rigorous evaluation protocol across all four settings. For the human baseline, participants have up to 30 minutes per question; if they cannot reach a reliable conclusion within this time limit, they may terminate the task and document their key exploration steps. For Tool-Augmented MLLMs, we enable the most advanced reasoning mode available to ensure unconstrained model performance. For Tool-Free MLLMs, models receive only the question text and original images without any tool access, and must directly generate the key information and final answer. Under the OmniSeeker setting, we limit interactions to a maximum of 20 rounds per question. The retrieval module uses Serper [^1] https://serper.dev and returns the top 5 results; image retrieval outputs are embedded into the dialogue context as base64-encoded data; the webpage access module uses Jina [^2] https://jina.ai to retrieve and parse webpage content; and image cropping is performed programmatically, with cropped images returned to the model.
Evaluation Metrics
We employ both result-level and process-level metrics. At the result level, we use Success Rate to measure whether tasks are completed successfully. At the process level, we introduce Process Score to quantify how much progress a model makes toward problem resolution during multi-step search and reasoning—specifically, the proportion of critical sub-goals successfully completed. This metric is formally defined as:
$\mathrm{ProcessScore}(q)=\frac{|\hat{\mathcal{G}}_{q}|}{|\mathcal{G}_{q}|},$ (1)
where $\mathcal{G}_{q}$ denotes the set of ground-truth sub-goals required to solve problem $q$, and $\hat{\mathcal{G}}_{q}$ denotes the set of sub-goals achieved by the model or human.
Table 2: Performance on BrowseComp-V3. Results are reported in terms of Success Rate and Process Score under the Pass@1 setting. Avg. denotes the average performance, while Sci., Tech., Soc., Cul., and Lif. correspond to the Science, Technology, Society, Culture, and Life categories, respectively. Bold numbers indicate the best-performing model within each group.
Model Success Rate (SR, %) Process Score (PS, %)
Avg. Sci. Tech. Soc. Cul. Lif. Avg. Sci. Tech. Soc. Cul. Lif.
Human
Browser 68.03 72.00 70.00 73.33 68.00 54.00 82.93 87.54 85.25 84.19 82.73 74.32
Tool-Augmented MLMs
GPT-5.2-Thinking 39.13 26.00 48.00 38.67 37.33 46.00 66.05 61.11 79.74 54.87 71.41 64.73
Gemini-3-Pro-Preview 22.90 18.00 16.00 21.33 24.00 34.00 62.43 62.02 73.78 48.70 66.33 62.50
Claude-Sonnet-4.5-Thinking 18.33 22.00 16.00 17.33 21.33 14.00 47.73 58.61 58.41 34.79 54.33 35.68
Tool-Free MLMs
GPT-5.2 6.00 0.00 14.00 4.00 5.33 8.00 25.02 17.50 44.50 19.22 25.25 21.43
o4-mini 7.33 0.00 16.00 2.67 6.67 14.00 29.08 29.48 46.12 17.91 29.04 28.44
GPT-4o 2.67 0.00 10.00 2.67 1.33 0.00 11.26 6.78 26.52 9.45 7.60 8.66
Gemini-3-Flash-Preview 12.00 8.00 18.00 12.00 10.67 12.00 40.76 38.98 61.94 31.89 39.72 36.60
Claude-Sonnet-4.5 4.00 4.00 6.00 2.67 5.33 2.00 25.74 33.06 47.56 15.97 24.53 13.06
Doubao-Seed-1.8 9.00 8.00 16.00 1.33 13.33 8.00 34.74 36.26 51.00 22.28 39.87 27.96
MiMo-V2-Flash 3.00 2.00 4.00 2.67 4.00 2.00 8.12 4.52 17.28 5.43 9.40 4.70
Qwen3-VL-235B-A22B-Instruct 3.33 4.00 4.00 4.00 2.67 2.00 20.52 26.34 35.38 11.69 19.64 14.39
Qwen3-VL-8B-Instruct 1.00 0.00 2.00 1.33 0.00 2.00 6.64 3.28 15.36 6.39 4.09 5.48
OmniSeeker (Ours)
GPT-5.2 36.00 50.00 28.00 33.33 33.33 38.00 57.70 67.23 55.40 49.34 60.49 58.81
o4-mini 26.00 22.00 24.00 25.33 34.67 20.00 44.66 43.70 52.11 40.73 48.72 37.94
GPT-4o 11.41 14.00 14.00 10.67 14.67 2.00 24.15 22.32 43.36 17.35 25.93 13.04
Gemini-3-Flash-Preview 23.67 32.00 24.00 18.67 25.33 20.00 47.37 50.84 68.35 40.15 43.68 39.27
Claude-Sonnet-4.5 22.67 32.00 20.00 21.33 24.00 16.00 54.17 60.94 64.25 45.29 56.73 46.78
Doubao-Seed-1.8 33.67 42.00 28.00 37.33 38.67 18.00 58.44 52.94 69.70 57.46 66.27 42.42
MiMo-V2-Flash 16.67 18.00 12.00 10.67 29.33 10.00 31.33 34.84 35.04 24.45 42.43 17.76
Qwen3-VL-235B-A22B-Instruct 14.33 16.00 14.00 14.67 17.33 8.00 26.68 28.56 35.93 22.00 28.36 20.05
Qwen3-VL-8B-Instruct 5.33 2.00 2.00 9.33 8.00 2.00 13.40 8.02 19.30 15.57 15.55 6.38
4.2 Main Results
Based on the experimental results in Table 2, we summarize our key findings as follows:
(1) Performance Gap and Benchmark Difficulty. Humans significantly outperform all models on BrowseComp-V3tasks, achieving an average Success Rate of $68.03\%$ and a Process Score of $82.93\%$. In contrast, no model achieves more than $40\%$ SR. This gap both highlights the limitations of current MLLMs in multimodal deep search tasks and validates the benchmark’s ability to capture real-world search complexity.
(2) Critical Role of Tool Augmentation. Without tool access, most models achieve only approximately $10\%$ SR. Tool augmentation yields substantial performance improvements, indicating that parameterized knowledge alone cannot adequately capture dynamic, cross-modal evidence chains on the open web. This highlights the importance of external retrieval and interactive capabilities for deep multimodal reasoning.
(3) Effectiveness and Generalizability of OmniSeeker. Empirical evidence confirms that OmniSeeker provides a unified and efficient tool-calling framework. When equipped with OmniSeeker, all models consistently achieve substantial improvements, reaching performance comparable to specialized proprietary systems.
(4) Value of Process-Level Evaluation. We observe a notable gap between PS and SR, with PS typically exceeding SR. This indicates that while models can complete individual sub-goals, they often fail to maintain logical consistency across long-sequence tasks. Therefore, fine-grained process-level evaluation is essential for identifying where and why models fail, thereby revealing their capability boundaries.
(5) Competitive Performance of Open-Source Models. While proprietary models (e.g., GPT-5.2 OpenAI [2025b]) remain the top performers, high-performance open-source models are rapidly closing the gap. Notably, Doubao-Seed-1.8 ByteDance Seed [2025] achieves $33.67\%$ SR when equipped with OmniSeeker. This demonstrates that high-quality open-source models possess strong capacity for complex reasoning and provide a promising path toward developing cost-effective, high-performance web browsing agents.
Table 3: PS across Different Models and Levels
Model L1 L2 L3
GPT-5.2 0.6176 0.5528 0.5792
Claude-Sonnet-4.5 0.5708 0.5353 0.5186
Doubao-Seed-1.8 0.6185 0.5652 0.5838
MiMo-V2-Flash 0.3776 0.2638 0.3420
Qwen3-VL-235B 0.3262 0.2308 0.2715
5 Further Analysis
5.1 Fine-grained Analysis
Task Level
As shown in Table 3, model performance declines substantially as task complexity increases from Level 1 to Levels 2 and 3. This reveals that while models can effectively perform unitary visual search, they face significant challenges in inter-region integration and inter-image relational reasoning.
Search Depth
As illustrated in Figure 3 (Left), SR for both humans and models decline with increasing search depth, yet exhibit distinct patterns. Human performance drops sharply with longer search paths, whereas model performance declines more gradually. This discrepancy implies that models leverage internalized parametric knowledge as a compensatory mechanism to mitigate the impact of search complexity.
Ability Boundaries
Figure 3 (Right) reveals distinct bottlenecks for humans and models. Human performance limitations are primarily in TextSearch, due to constraints in attention span and cognitive load when processing voluminous text. In contrast, multimodal integration remains the primary bottleneck for all models.
Figure 3: Difficulty and Ability Analysis
Figure 4: Test Time Scaling
Figure 5: Failure Mode Analysis
5.2 Test Time Scaling
We evaluate how test-time compute affects performance on BrowseComp-V3. Our key findings are as follows:
  • Scaling Interaction Steps. As shown in Figure 4 (Left), increasing the maximum number of interaction turns substantially improves performance. Notably, Qwen3-VL-235B exhibits a stronger scaling advantage than its 8B counterpart. This suggests that larger models have stronger long-horizon reasoning capabilities, allowing them to better utilize additional interaction steps for iterative refinement.
  • Scaling Sampling Consistency. Figure 4 (Right) shows the performance of Qwen3-VL-235B as we increase the number of independent samples ($N$). Among the three strategies, Best-of-$N$ scales most effectively, continuously improving performance with increasing $N$.
5.3 Failure Mode Analysis
We analyze the error distributions of four representative models, as shown in Figure 5. Our key findings are summarized below:
  • Multimodal Grounding and Perception. Across all models, Visual Grounding and Perception Failure dominate the error distribution. This indicates that current MLLMs struggle to accurately extract and perceive visual information in complex, noisy web environments.
  • Multimodal Progress, Planning Constraint. Closed-source models substantially reduce perception and grounding errors compared to open-source models. However, with improved multimodal capabilities, long-horizon planning becomes the main bottleneck limiting further improvements in SOTA models.
6 Conclusion
In this work, we introduce BrowseComp-V3, a comprehensive benchmark for the evaluation of multimodal deep browsing and search capabilities. The benchmark consists of 300 rigorously curated and annotated questions designed to systematically remedy three core limitations of existing evaluation paradigms: task complexity, information searchability, and evaluation dimensions. Empirical results reveal that SOTA MLLMs achieve under 40% SR, underscoring a substantial gap relative to human performance. These findings confirm the effectiveness and discriminative power of BrowseComp-V3in simulating open-world multimodal deep search scenarios. Further analysis reveals critical deficiencies in current models’ capacity to integrate and comprehend multimodal information, whereas the process-level evaluation and Test-Time Scaling analysis offer potential pathways for enhancing model capabilities via methodologies such as reinforcement learning. Additionally, our agent framework, OmniSeeker, achieves performance comparable to leading closed-source models, offering an open alternative for developing multimodal browsing agents. In conclusion, BrowseComp-V3provides a comprehensive platform for evaluating and advancing multimodal browsing agents. Its process-level evaluation and fine-grained capability analysis will catalyze future breakthroughs in multimodal deep search.